### v3

# Decentralization of Ethereum’s Builder Market

Abstract
Blockchains protect an ecosystem worth more than \$500bn with strong security properties derived from the principle of decentralization. Is today’s blockchain decentralized? In this paper, we empirically studied one of the least decentralized parts of Ethereum, its builder market.
The builder market was introduced to fairly distribute Maximal Extractable Values (MEV) among validators and avoid validator centralization. As of the time of writing, three builders produced the vast majority (more than 80%) of blocks in Ethereum, creating a concerning centralization factor. However, the community believes that such centralization “is okay,” arguing that builder centralization will not lead to validator centralization. In this empirical study, we interrogate the causes and implications of builder centralization and challenge this belief that it is acceptable.
Our main finding is that builder centralization has led to a significant loss by validators and, if left uncontrolled, could lead to validator centralization. Moreover, MEV mitigation solutions slated for adoption are affected too because they rely on the builder market as an “MEV oracle,” which is made inaccurate by centralization.
Our investigation revealed two reasons behind builder centralization. We propose a structural change to the existing MEV supply chain and a solution based on the new supply chain structure. However, future work is required to analyze if the new supply chain structure is sustainable in the long term, which we leave open.
1 Introduction
At the core of any decentralized blockchain, a key assumption underpinning their strong security properties (integrity, availability, censorship resistance, incentive compatibility, etc.) is “decentralization”, that the system is run by a large number of independent participants [63, 28, 75, 83, 4, 74]. It is community consensus that one of the greatest threats to decentralization is Maximal Extractable Value (MEV) [16], profits that protocol participants, known as validators, can reap by manipulating the ordering of transactions. MEV can lead to centralization because it disproportionally benefits well-resourced validators compared to “regular” ones, as extracting MEV requires significant capital, intelligence, and computational resources, which regular validators may not have [4, 23, 67, 12, 11].
To distribute MEV fairly among validators and avoid validator centralization, Ethereum adopted a solution called Proposer-Builder Separation (PBS) [23]. A proposer is the validator chosen to add a new block to the blockchain. In the current implementation called MEV-Boost [35], the proposer runs an auction among builders (anyone can be a builder) to solicit the best block. Ideally, builders compete with each other, and the proposer collects a significant portion of MEV in auction revenue without doing much work, eliminating the advantage of sophisticated validators over regular ones.
Figure 1: Market share of builders. The builder market is one of the least decentralized parts of Ethereum!
Is the builder market working as intended? The answer is rather unclear because the builder market is highly centralized. As shown in Fig. 1, three builders won over 90% of auctions, raising questions about the competitiveness of the market. However, the community believes that “it is OK to centralize block building” [23], arguing that block centralization will not lead to validator centralization. We challenge this belief through an empirical study of the past MEV-Boost auctions. As will be shown in detail, builder centralization has caused a significant loss to the proposers and, if left uncontrolled, could lead to undesired consequences including validator centralization.
1.1 Implications of Builder Centralization
Proposer loss and its consequences.   Intuitively, a proposer’s loss is the difference between the MEV available in a block and the auction revenue. Builder centralization can cause proposer loss in two ways. First, auctions may not be competitive enough due to, e.g., collusion or other limitations of the auction mechanism. Second, even if builders compete with each other, if a single builder has a unique advantage, it will not be incentivized to bid on all of the MEV it can extract. Instead, it only needs to outbid the second-best bidder slightly to win. We say that the proposer incurs a loss from uncompetitiveness in the first case and a loss from inequality in the second case.
If proposers suffer significant losses, they would be incentivized to extract MEV by themselves (instead of using the builder market). This is undesirable because it re-introduces the dynamics that PBS wants to avoid: sophisticated validators would have an advantage over small ones in extracting MEV. In other words, significant proposer losses would make PBS “unstable”. Moreover, several MEV mitigation solutions (e.g., MEV-burn [19]) rely on the MEV-Boost auctions as an “oracle” to know how much MEV has been extracted from the system (e.g., MEV-burn proposed to remove that amount from circulation, effectively redistributing them to all holders of ETH through deflation.) The proposer’s loss translates to inaccuracy in the MEV oracle.
Next, we introduce the auction mechanism and present our method to quantify the proposer losses using auction data.
Modeling MEV-Boost auctions.   MEV-Boost auction adopts an open-bid, ascending price auction akin to an English auction [6], with a 12-second deadline. For each auction instance, builders submit bids in the form of $(B,BV)$ specifying a block $B$ and a bid value (BV) the builder is willing to pay for $B$ to be added to the blockchain. The revenue a bidder gains from $B$ is called the true value ($TV(B)$), which can be estimated from transactions in $B$. At the end of an auction, the block $B$ with the highest bid is added to the blockchain, the builder pays $BV$ to the proposer and earns $TV$ when $B$ is executed (i.e., the builder’s profit is $TV-BV$).
Recall that the proposer’s loss is the difference between the auction revenue and the MEV available in a block. The latter can be represented by the builders’ highest true values. With that, the two types of losses can be computed from auction data, as illustrated in Fig. 2. Specifically, suppose the true values of the bids from a given auction instance are $TV_{1}>TV_{2}\geq\cdots\geq TV_{N}$. The loss from inequality is $TV_{2}-TV_{1}$ because a competitive auction can, at most, generate revenue of $TV_{2}$. The loss from uncompetitiveness is $\max(TV_{2}-BV,0)$ (if this value is negative, the auction is considered competitive, and there is no loss from uncompetitiveness.)
Figure 2: Definition of proposer losses.
Auction data.   A practical challenge in quantifying the proposer loss in past MEV-Boost auctions is the lack of data. Only the winning bids are recorded in the blockchain, but to compute the losses, we also need losing bids. The key ingredient that enabled our study is a large-scale auction dataset we curated since 2022: we collected 8.79 billion partial bids [^1] Partial and full bids are to be defined in Sec. 3.1 of auctions from September 2022 to July 2024; in addition, in collaboration with ultra sound relay [62], we obtained full bids for a total of 301,479 auctions from April 2023 to December 2023. From the auction dataset, we can compute the true value of all bids.
Results.   91.8% of the MEV-Boost auctions we analyzed were competitive. However, auctions tend to be less competitive when the MEV of a slot increases. The loss from uncompetitiveness turns out rather insignificant, ranging from 0.5% to 1.7% during the periods of our study.
To understand the variations in builders’ abilities to extract values, we statistically compared the true values across different builders in the same auction. We used several metrics, including the quartile coefficient of dispersion (QCD) [88], to quantify the inequality of true values across builders. We found significant differences in the block-building capacities of top, middle, and tail builders. The inequality worsens as the MEV of a slot increases, likely because significant MEV opportunities are shared with only a limited number of builders.
Inequality of block-building capacities has led to a proposer loss ranging from 5.6% to 11.5%. I.e., had there been no inequality, the proposer would have earned roughly 10% more. This indicates that inequality is the primary driver of the proposer’s losses in the MEV-Boost auctions.
1.2 What caused Inequality in Block-building Capacities
Given that proposers suffer significant losses due to the builders’ block-building capacity inequality, we now investigate the underlying causes of this inequality.
In the current MEV ecosystem (to be presented in Sec. 2.2), builders receive order flows from order flow providers and compute the optimal ordering of received order flows (see Appx. E for a reference builder algorithm). Thus, a builder’s block-building capacity is determined by the “quality” of the order flows it receives. At the same time, not all order flows are accessible to all builders. The difference in builder’s ability to access important private order flows drives the inequality in block-building capacity. This also explains builder centralization: only a small set of builders can access these profitable order flows.
We proposed a metric to identify important order flows, and we investigated their accessibility by different builders and the underlying reasons.
Identifying pivotal private order flow providers.   To formally capture the intuition of “important providers”, we defined pivotal level to quantify the sustained impact of a given provider. A provider $P$ is pivotal in an auction $A$ if the winner of $A$ would have lost, had transactions from $P$ were not available. In other words, pivotal providers are necessary for the winner to win. The pivotal level of a provider $P$ is the percentage of MEV-Boost auctions in which $P$ is pivotal.
Using the auction dataset, we identified five pivotal providers whose pivotal level exceeds 50% over a period longer than two weeks. That is, if a builder cannot access the flows from any of these providers, it will lose the majority of the auctions during our study period.
These order flows are not equally inaccessible to all builders for two reasons we now introduce.
Inequality due to trust barriers.   Order flow providers do not openly share private order flows with any builder because a malicious builder can harm the provider by, e.g., sandwich attacks [94] and imitation attacks [71]. Providers set up different access barriers for self-protection.
Most MEV searchers (a specific type of order flow providers) resort to “obscurity for security” in that they are not publicly reachable; thus, we could not learn about their access barriers and how to overcome them. On the other hand, MEV channels (another type of provider) typically publish a clear access policy, which allowed us to gain some insight into the nature of their access barrier.
For example, MEV-Blocker only shares order flows with builders with a market share of at least 1%, using market share as a proxy for reputation and trustworthiness. This requirement introduces a chicken-and-egg problem for new builders: to gain market share, they need access to private order flows to win auctions, but as newcomers, they lack the required market share to access these flows. As a result, new builders often resort to subsidization: bidding higher than their actual profits in MEV-Boost auctions to establish the required market shares. Therefore, a new builder needs to keep subsidizing until reaching a certain market share to overcome the access barriers set by MEV channels. The cost to do so continues to increase over time and can reach up to 7.4 ETH, according to our quantification.
In summary, with a moderate cost, builders can overcome the access barrier of one type of order flow provider (MEV channels). However, how to access order flows from other providers, particularly important searchers, remains unclear.
Inequality due to integration.   A more worrisome finding is that some providers send order flows almost exclusively to their preferred builders, causing an inequality in block-building capacities.
To identify such provider-builder pairs, we introduced builder-specific pivotal level of a provider $P$ for a builder $B$ as the pivotal level of $P$ in auctions won by $B$. With this metric, we rediscovered known exclusive relationships (called “integration” [40]), but also found unreported integrated providers for top builders (Titan and Beaver respectively). This finding is interesting because Titan has been the only top builder without a well-known integrated provider. Our findings show that all three top builders, accounting for 90% of market share, have integrated providers.
Integration poses a market entry barrier as new builders likely lack integrated providers. Moreover, it offers a unique competitive edge and will likely cause dominance. Finally, there is a strong incentive to integrate: compared to sending valuable order flows to multiple builders, sending them to a single builder enables this builder to win the auction with higher profit; the provider and the builder can share the profit, which means integration improves their total utility.
Integration is detrimental to proposers, as it will reduce their profits. The strong incentive for integration makes it increasingly difficult to address builder centralization.
1.3 Mitigation Ideas
To summarize, we found that trust barriers and incentives for integration contribute to builder centralization, leading proposers to incur significant losses. The consequences include the instability of PBS and the inaccuracy of MEV oracles.
We propose a structural change to the existing MEV supply chain (Sec. 2.2) and a solution based on the new supply chain structure. However, future work is required to analyze if the new supply chain structure is sustainable in the long term, which we leave open.
The new supply chain has only one builder to which the order flow providers send flows and from which the proposers accept blocks. In this setting, a recent work [57] suggests a truthful mechanism that incentivizes order flow providers to report the MEV in their order flows truthfully.
Our solution involves having the builder run the truthful mechanism in [57] in a Trusted Execution Environment (TEE) [15], hardware-based enclaves that guarantee the confidentiality and integrity of programs running inside. Order flow providers verify that the correct mechanism is running in the TEE using remote attestation, then send order flows to the TEE along with a bid indicating the MEV in their order flow. Then, the TEE computes a fair refund to order flow providers. The refund mechanism is designed so that bidding truthfully is a dominant strategy [57]. (While oversimplifying, one can make a loose analogy to sealed-bid second-price auctions among order flow providers.) Then, the TEE can compute the total MEV in a block from the providers’ bids, solving the problem of inaccurate MEV oracles.
This also addresses the instability of PBS because the builder program in the TEE can be programmed to pay the proposer fairly, leaving just enough for its operation. Since this builder is fair to order flow providers, it is conceivable that it will be preferred by order flow providers over other builders. However, as mentioned above, it is unclear if having a single builder is an equilibrium. Also, the proposers may be incentivized to become order flow providers, which may re-introduce the concern similar to PBS instability. We elaborate on the idea in Sec. 7.1.
Contribution
To the best of our knowledge, we are the first to use the rich auction data to analyze the decentralization of Ethereum’s builder market, while most previous works are limited to on-chain data. We summarize our contribution below.
  • We found that builder centralization has led to significant proposer losses, which could lead to undesirable consequences, including validator centralization and inaccurate MEV oracles. We challenge the community’s belief that “builder centralization is okay".
  • We attribute proposer losses to two factors: auctions are not competitive, builders’ block-building capacities are unequal, and inequality is the primary cause.
  • We identify two reasons for inequality in block-building capacity: providers’ trust concerns and the incentive for integration.
  • Our dataset of the true values of 301,479 MEV-Boost auctions from April to December 2023 may be of independent interest. Our dataset of 8.79 billion partial bids since 2022 is the largest scale known to us.
  • Besides contributing to the literature on MEV mitigation, our work also contributes to the Economics literature on the analysis of MEV-Boost auctions. Currently, there is only limited auction theoretic analysis that considers the complexity of real-world MEV auctions [40, 67, 90]. The findings in this paper can serve as empirical evidence for future theoretical interrogation.
Figure 3: Illustration of the MEV supply chain in the PBS scheme.
2 Background
2.1 MEV-Boost Auctions
MEV.   Miner/Maximal Extractable Value (MEV) [16] refers to the profit that privileged players (e.g., validators) can earn by including, excluding, and reordering the transactions. MEV poses a threat to decentralization, as explained in the introduction. There are two schools of thought on mitigating the negative impacts of MEV. The first one is referred to as MEV democratization [30], the main idea of which is to facilitate MEV extraction and level the playground. MEV-Boost [35], to be introduced below, is the primary example. Another approach is to enforce certain ordering policies (e.g., FIFO) to prevent order manipulation, such as [45, 44]. We focus on MEV-Boost because it is widely used.
PoS Ethereum.   Before introducing background on MEV-Boost auctions, we first cover the relevant concepts in Ethereum’s Proof of Stake-based consensus protocol called Gasper [13]. Participants of Gasper are called validators. Each validator must deposit 32 ETH as collateral [22]. Time is divided into 12-second intervals called slots. This results in 7,200 slots in Ethereum per day. During each slot, a single validator, referred to as the proposer, proposes a block, while other validators vote for the chain head in their local view.
Proposer-builder separation (PBS) and MEV-Boost.   PBS can be conceptually viewed as a sidecar to the core consensus protocol (although some proposals are more tightly coupled with consensus) to allow the proposer to outsource the job of building blocks to builders.
MEV-Boost [35] is the current incarnation of the PBS idea. About 90% of all blocks on Ethereum are built by MEV-Boost [87]. The basic idea is to auction off the right to propose a block to the builder who bids the highest in an open-bid ascending price auction [14]. Because builders and proposers are mutually untrusted, the auction is facilitated by a (trusted) third party called a relay [29]. Here are the high-level steps (we refer readers to [35] for details):
  • Builders submit bids to relays throughout the slot. For this paper, we denote a bid with $(B,BV,PK_{\text{builder}})$ that consists of a block $B$, a bid value $BV$ to be paid to the proposed if the block is proposed, and the builder’s public key. Bids are signed by the builder but we omit the signature for clarity. Relays verify that $B$ indeed pays the proposer at least $BV$.
  • Relays expose a public API that publishes partial bids in the form of $(\mathsf{Hash}(B),BV,metadata)$. $metadata$ includes parent hash, gas limit, gas used, etc. Partial bids do not include block content.
  • The proposer selects a block hash $h_{B}$, typically the one with the highest $BV$, and sends the relay a signature on $h_{B}$. This signature is a promise to propose $B$ in the consensus protocol (if it proposes another block, this signature is evidence of equivocation and will cause the proposer to be penalized [13].) After verifying the signature, the relay returns $B$ to the proposer.
Two additional remarks. First, relays are trusted entities, but anyone can run a relay (currently ten [27]), and builders and proposers can connect to the relays they trust. Second, winning bids are recorded on-chain, and losing bids are discarded after the auction. We collected and archived the historical bids over time to enable this study, as we will detail in Sec. 3.
2.2 MEV Supply Chain
The ecosystem around discovering and extracting MEV forms an MEV supply chain, as depicted in Fig. 3. Builders, relays, and proposers (validators) have been introduced above. Now we introduce the “upstream” of builders.
Wallet, users, trading bots.   Users use wallets to interact with blockchains. A user communicates her intent (e.g., transferring cryptocurrency) to the wallet, which generates the transaction to fulfill the user’s intent and sends it to the network. MEV may be generated during this process, e.g., when the user transaction creates an arbitrage opportunity. While the wallet is typically controlled by the users, it can also be managed by a third party, such as Telegram trading bots [43]. These bots are automated programs integrated with the Telegram app [79], automatically conducting cryptocurrency trades on behalf of users. Users’ transactions eventually reach builders or proposers, either through the peer-to-peer network (the mempool) or through third-party services called private channels.
Mempool.   Mempool is a temporary storage of pending transactions before they are added to the blockchain. All transactions stored in the mempool are publicly visible.
Channel.   A private channel, or channel for short, connects users directly with builders, bypassing the mempool. Channels receive transactions from users and then, forward them to the user-specified builders, offering privacy and atomicity guarantees. E.g., a user can send a bundle (list) of transactions and request them not to be unbundled. Examples include MEV-Share [37], MEV Blocker [60], BloXroute [9], etc.
Searcher.   Searchers are a special type of users. They run (typically proprietary) algorithms to construct profitable transaction sequences (bundles) based on the pending transactions and the blockchain state. Same as regular users, searchers may submit transactions to mempool or channels. In addition, most builders have public APIs that allow searchers to submit bundles directly. Some searchers may prefer channels for the privacy and atomicity benefits. A searcher is said to integrate with a builder if it sends almost all of its orders to that builder exclusively. For example, searcher 0xA69b… is considered an integrated searcher of Beaver Builder because Beaver Builder includes the majority of its transactions.
Some channels, such as MEV-Share and MEV Blocker, allow searchers to read pending transactions in the channel and extract MEV from them, provided that such extraction should not harm users (e.g., only backrunning is allowed).
Private order flow.   Order flow refers to a stream of transactions that may carry extractable value and can be categorized into two categories: public order flow from the public mempool, and private order flow, sent directly by users or searchers without being broadcast to the mempool. The entities that can provide private order flows to the builders are referred to as private order flow providers, which include searchers, channels, users, and Telegram trading bots as shown in Fig. 3.
2.3 Model and Assumptions.
Accounts and states.   We follow the model of [2, 3]. We use $\mathcal{A}$ to denote the space of all possible accounts. For example, in Ethereum, both user-owned and contract-owned accounts are represented by 160-bit identifiers. For $a\in\mathcal{A}$, $\text{balance}(a)$ denotes the balance of all tokens held in account $a$, and $\text{balance}(a)[\mathbf{T}]$ specifies the balance of token $\mathbf{T}$. For simplicity, $\text{balance}(a)[0]$ is used to denote the balance of the primary token (e.g., ETH in Ethereum).
Transaction and block.   A transaction $tx$ is a string whose execution changes the system’s state. A block is an ordered sequence of transactions. Executing transactions in a block $B$ transforms an initial state $s$ into a new state $s^{\prime}$, represented by $s^{\prime}=\text{action}(B)(s)$.
Extracted value (EV).   The extracted value of a block $B$ in state $s$ for a participant $p$ is the changes of $p$’s balances after executing $B$. Assuming $p$ controls a set of accounts $A_{p}$, the extracted value of $B$ for $p$ is:
$EV_{p}(B)=\sum_{a\in A_{p}}\text{balance}_{s^{\prime}}(a)[0]-\text{balance}_{s}(a)[0]$ (1)
where $s^{\prime}=action(B)(s)$.
Builder’s income.   We assume a builder $p$’s income from building a block $B$ is $EV_{p}(B)$, or $EV(B)$ for short since the builder identity is usually encoded in the block. We compute EV following a similar approach as that in MEV analysis tools [20, 31] and previous works [41, 80]. Specifically, $EV(B)$ can be computed as the sum of the transaction priority fees [21] and direct transfers [32] to the builder’s addresses minus the refund to other addresses. We also refer to the builder’s income from a block (or a set of transactions in a block) as the value of the block (or the set of transactions).
Note that $EV(B)$ is a lower bound of the builder’s actual income from $B$ because the builder may own accounts out of $A_{p}$ in Equation 1, or even receive payments off-chain. We discuss the implications in Sec. 7.
3 Dataset
We created two auction datasets to study the decentralization of the builder market: a dataset of partial bids from September 2022 to July 2024, and a dataset of full bids from April to December 2023. The full bids are obtained from the ultra sound relay [62]. Recall from Sec. 2.1 that a bid consists of a block $B$ and a bid value $BV$ and a partial bid consists of $(\mathsf{Hash}(B),BV,PK_{\text{builder}},metadata)$ where $metadata$ includes parent hash, gas limit, gas used, etc. To support analysis, we also collected on-chain data, including blocks and transactions on Ethereum, as well as miscellaneous data, which will be detailed below. Tab. 1 summarizes the datasets.
Dataset Content # of entries Source
On-chain blocks 4,892,579 Reth node
transactions 748,859,795
Auction blocks produced by MEV-Boost 4,305,571 relay API
partial bids 8,789,557,915 relay API
full bids from ultra sound 326,574,676 ultra sound relay
Misc. builder public keys 248 see Sec. 3.1
builder addresses 102
searcher addresses 2,140
private transactions 77,747,623
transaction sources 60,345,066
Table 1: Overview of collected datasets.
3.1 Data Collection and Validation
On-chain data.   We run an Ethereum execution client Reth [68] along with a consensus client Lighthouse [69] to collect blocks and transactions from September 15, 2022, to July 31, 2024. The on-chain dataset includes 4,892,579 blocks and 748,859,795 transactions.
Data from relay APIs.   We connected to all 13 known relays [^2] Three relays were inactive during the period of our study. and collected the following information from their public APIs: 1) the winning bid including the final block sent to the proposer; and 2) partial bids for blocks that did not win the auction. Recall that partial bids do not include the block body. We collected 4,305,571 winning bids and 8,789,557,915 partial bids from September 15, 2022, to July 31, 2024. We use winning bids to identify blocks produced through MEV-Boost. We use partial bids to quantify the importance of private order flow providers in Sec. 5.
To validate the completeness of the data collected from relay APIs, we compared the winning bids we collected with three public datasets [64, 85, 17]. Our dataset includes 720 more winning bids than other datasets combined. Upon comparison with on-chain blocks, we find they are all included on Ethereum, thereby indicating a high level of completeness.
Full bids from the ultra sound relay .   From the ultra sound relay [62], we obtained the full bids in 301,479 historical auctions that took place between April and December 2023 (ultra sound relay stopped collecting bids at the end of 2023). Since the data volume is large (we record roughly 200 GB of data per day), our dataset samples one week from each month: April 9-15 and May 1-7 to December 1-7, 2023. The dataset comprises 326,574,676 full bids (block, bid value, and builder public key). We use full bids in Sec. 4 to quantify the competitiveness and efficiency of MEV-Boost auctions.
The original data from relays includes bids submitted after the deadline because of network latency. We removed them as we are only interested in the bids before the auction ends. On April 18, 2023, MEV-Boost introduced bid cancellation that allows a bidder to cancel a previous bid by submitting a lower one. However, proposers can query bids before they are canceled and store them locally [24], effectively bypassing cancellation. Therefore we do not consider cancellation.
Not all builders connect to the ultra sound relay, so we need to validate that the full bids from ultra sound relay cover a significant subset of all builders. We compute the percentage of the builders that submit to ultra sound relay over all known builders. The set of all known builders is a union of all builders that appeared in our datasets, including winning bids, partial bids, and bids from the ultra sound relay. As shown in Fig. 10, bids from the ultra sound relay covered more than 80% of the builders in over 75% of the MEV-Boost auctions during our study period. Moreover, the top-5 builders by market share (Flashbots, Beaver, Rsync, Builder0x69, and Titan) almost always submitted to ultra sound relay.
Miscellaneous data.   In practice, builders are known by their common names (e.g., flashbots builder), but a builder may use multiple public keys to sign bids and use multiple addresses to receive and make payments. We collect builders’ public keys to attribute bids to builders. We collect builder’s addresses to calculate the true value of bids. We use private transaction hashes to identify private order flows and use the searchers’ addresses and sources of transactions to identify the private order flow providers in Sec. 5.
  • Builder public keys: We collect public keys controlled by each builder combining two data sources: the public keys in builders’ official documents; and the distinct marks in their blocks’ “extra field” (see Appx. B for details).
    To validate the completeness of our dataset, we manually cross-check it with public datasets [72, 65] and related works [42, 41], and the results show that our dataset includes 119 more public keys, so our dataset is the most complete one so far.
  • Builder addresses: We collect the builder addresses from the blocks’ last transactions combined with the public dataset [26]. Similarly, to validate their completeness, we manually cross-check them with related works [42, 41], and the results show that our dataset includes 86 more builder addresses, which is the most complete so far.
  • Searcher addresses: We obtained known searcher addresses from Etherscan [25], libMEV [50], and related work [42].
  • Private transaction hashes: We use data from Mempool Guru [92] to label transactions that bypassed the public mempool as private. We believe these labels have already been cross-validated since the private transactions are collected by seven distributed nodes.
  • Transaction sources: we are interested in where a transaction originates from and which part of the MEV supply chain it has traversed through. We identified transaction sources using ground truth data (as described in Appx. B).
4 Implications of Builder Centralization
In this section, we focus on the status quo and study the implications of a centralized builder market as we have today. In particular, we are interested in the losses incurred by proposers, as significant losses could lead to undesirable consequences, including validator centralization and inaccurate MEV oracles.
We quantify proposers’ losses in historical MEV-Boost auctions considering two factors: the competitiveness of the auctions and the inequality in block-building capacity.
Our study includes some MEV-Boost auctions where the highest bids were not included on Ethereum due to network latency. However, we do not omit these auctions since the auctions themselves remain meaningful.
4.1 Competitiveness of MEV-Boost Auctions
Ideally, revenue from MEV-Boost auctions should be near optimal due to competition. However, whether real-world auctions induce such near-perfect competition is complicated by several confounding factors: builders may not be able to respond to bids due to network latency, or they may collude implicitly or explicitly. In this section, we quantify the competitiveness of the MEV-Boost auctions using what we call competitive index (CI).
We also look at the allocation efficiency as it evaluates whether the current mechanism induces the desired outcomes.
True value.   In the context of auctions generally, the true value of the item being auctioned refers to its actual worth to the bidder. Recall that in MEV-Boost auctions (Sec. 2), a builder submits a block $B$ (a list of transactions) and a bid value $BV$ that the builder is willing to pay to the proposer if $B$ is proposed. The true value of $B$, denoted $TV(B)$, is typically higher than $BV$. One can think of $TV$ as the builder’s revenue and $BV$ as the cost. The difference $EV:=TV-BV$ is the profit pocketed by the builder, which we call extracted value (EV) — i.e., $EV(B)$ is the difference of the builder’s ETH balance after executing transactions in $B$. For example, suppose executing $B$ yields a net profit of $0.1$ ETH after paying $0.9$ ETH in bid value in the auction, then the true value (the revenue) is $1.0$ ETH.
To compute $TV(B)$ of a given block $B$, we follow the standard approach in open source MEV analysis tools [20, 31] and previous works [41, 80], though we built our own analysis toolchain and optimized its performance to handle the large volume of auction data. We compute $EV(B)$ using a modified Ethereum execution client (Reth [68]) to execute transactions in $B$ and compute the net ETH balance changes of the bidder’s addresses. Most builders’ addresses are well-known partly because they need to build up a reputation. We cross-validated it with other sources to ensure that our dataset was the most complete one to date (see Sec. 3.1 for details.) Then, with bid value $BV(B)$ from the auction dataset, the true value $TV(B)$ is the sum of $EV(B)$ and $BV(B)$.
In every slot $s$, bidders can submit multiple bids throughout the duration of the slot. We use $BV(p,s)$ to denote the highest bid value from bidder $p$ in slot $s$, and $TV(p,s)$ the corresponding true value.
Competitive index.   In a given slot $s$, suppose builders $p_{1},\cdots,p_{n}$ are ordered by their true value from high to low, i.e., $TV(p_{1},s)\geq\dots\geq TV(p_{n},s)$, $BV_{w}(s)$ is the bid value of the winner and $TV_{w}(s)$ is the corresponding true value, the competitive index of slot $s$, $CI(s)$ is defined as:
$CI(s)=\frac{BV_{w}(s)-TV(p_{2},s)}{TV(p_{2},s)}\times 100\%.$
Namely, $CI(s)$ measures the relative difference between the winning bid and the second-highest true value. $CI(s)\geq 0$ indicates that the winning bid value is not less than the second-highest true value, satisfying the definition of a competitive auction, while $CI(s)<0$ means insufficient competition.
Efficient index.   Same as above, the efficient index, $EI(s)$, is defined as:
$EI(s)=\frac{TV_{w}(s)-TV(p_{2},s)}{TV(p_{2},s)}\times 100\%.$
$EI(s)>0$ indicates that the winner has the highest true value, whereas $EI(s)\leq 0$ implies that the bidder with the highest true value lost the auction.
Figure 4: The distribution of CI and EI, where each dot represents an auction. The color represents density when multiple dots overlap. 91.8% of the MEV-Boost auctions are competitive (sum of the slots having $CI\geq 0$), and 75.1% of them are efficient (sum of the slots with $EI>0$). Dots on the diagonal line ($CI=EI$) are auctions where the winners bid the true values. The dots above this line are auctions where the winners bid lower than true values.
Analysis of uncompetitiveness and inefficiency.   Fig. 4 shows the calculated CI and EI distribution across all MEV-Boost auctions in the ultra sound relay dataset. Notably, 91.8% (sum of the areas with $CI\geq 0$) of the auctions are competitive, yet only 75.1% (sum of the areas with $EI>0$) are efficient.
We note Fig. 4 shows a somewhat strict definition of competitiveness (i.e., $CI\geq 0$). In reality, builders may bid slightly lower than their true value to ensure profits. Therefore, it is reasonable to relax and consider an auction with $CI\geq-\delta$ to be competitive for a small $\delta$. We call this notion $\delta$-competitive. As shown in Fig. 13, when we consider $\delta=1\%$ or $\delta=2\%$, the percentage of uncompetitive auctions reduces to about 6% and 5%, respectively. That is, under the slightly relaxed definition, approximately 95% auctions were competitive.
In the interest of space, we leave the analysis of efficiency and the implication of MEV on efficiency to Appx. D.1.
MEV’s implication on auction outcomes.   To further understand how the magnitude of MEV in a slot affects the competitiveness and efficiency of the auction, we categorize every slot into one of three tiers according to the winning bid (as a proxy to the significance of the MEV opportunity in that slot): low MEV (below 0.0215 ETH, the 10th percentile of the winning bid values), high MEV (above 0.3933 ETH, 95th percentile) and medium MEV in between. The ratio of competitive auctions in each tier is shown in Fig. 5. We note that competitiveness decreases as the MEV of a slot increases.
Figure 5: The ratio of auctions that are competitive and efficient across various MEV tiers.
Finding I: MEV-Boost auctions are generally competitive, while the competitiveness of MEV-Boost auctions deteriorates as a slot’s MEV increases.
Loss from uncompetiveness.   For an uncompetitive auction, the loss from uncompetitiveness is the difference between the second-highest true value and the actual winning bid value, representing the loss a validator incurs compared to what it would have received in a competitive auction. We computed the loss from uncompetitiveness using our dataset, and the results are shown in Tab. 2. The loss from uncompetitiveness ranges from 0.5% to 1.7% during the periods covered.
Table 2: Losses from uncompetitiveness and inequality.
Time Slots Profits (ETH) Losses-un (ETH) (%) Losses-in (ETH) (%)
April 9-15 28,385 2,704.4 46.9 (1.7) 312.1 (11.5)
May 1-7 30,300 9,331.7 115.8 (1.2) 518.6 (5.6)
June 1-7 35,443 4,341.8 25.1 (0.6) 342.2 (7.9)
July 1-7 36,040 3,938.8 19.1 (0.5) 246.1 (6.3)
August 1-7 17,831 2,135.5 12.5 (0.6) 146.6 (6.9)
September 1-7 35,815 3,939.5 48.5 (1.2) 247.7 (6.3)
October 1-7 39,131 3,427.0 28.9 (0.8) 234.1 (6.8)
November 1-7 39,848 4,736.2 45.6 (1.0) 305.0 (6.4)
December 1-7 38,686 5,884.4 80.0 (1.4) 365.8 (6.2)
4.2 Inequality in Block-building Capacities
The previous section shows that the MEV-Boost auctions are competitive in most cases. However, a proposer may still incur loss due to a significant block-building capacity gap between builders. In MEV Boost auctions, this quantity is captured by the true values of the builder’s bids.
Quantifying the inequality of block-building capacities.   We compare the highest true values of different builders in the same slot to understand the differences in builders’ ability to extract values. If a builder consistently has the highest true value, then she can potentially dominate the builder market. On the contrary, if many builders have similar true values, then it is plausible that a properly designed market could avoid a monopoly.
To quantify the disparity of block-building capacity, we compute the quartile coefficient of dispersion (QCD) [88] of the highest true values in a given slot. Specifically, suppose the true values of different builders in slot $s$ are $\vec{tv}_{s}$. Let $Q_{1}(\vec{x})$ and $Q_{3}(\vec{x})$ representing the first and third quartiles of $\vec{x}$ (i.e., 25% and 75% percentiles). The QCD of slot $s$ is
$QCD_{s}=QCD(\vec{tv}_{s})=\frac{Q_{3}(\vec{tv}_{s})-Q_{1}(\vec{tv}_{s})}{Q_{3}(\vec{tv}_{s})+Q_{1}(\vec{tv}_{s})}.$
QCD is a real value in $[0,1]$. A larger QCD indicates greater dispersion (the intuition is that in a dataset with big dispersion, $Q_{3}$ is much larger than $Q_{1}$, thus $QCD$ approaches $1$). Since it uses quartiles, QCD is robust to outliers [88].
We now present the result given by QCD, noting that other metrics lead to a similar conclusion (see Sec. D.2).
Evaluation results.   We categorize the builders into three groups based on their market share from April to December 2023: top builders (top 5), middle builders (6-15), and tail builders (16-25). To understand how the magnitude of MEV in a slot affects the block-building capability, we follow the same methodology in Sec. 4.1 and categorize the auctions into three tiers: low MEV, medium MEV, and high MEV.
In Fig. 6, we plot the distribution of the QCD for MEV-Boost auctions from April 2023 to December 2023 for different builder groups and MEV tiers. (We sample the first week of each month. See Sec. 3.1 for details.) It reveals two trends. First, within each MEV tier, there is a clear increase in inequality (a high QCD means high inequality) from the top to the middle and then to the tail. This suggests that while top builders have comparable block-building capacities, middle and tail builders have significantly different capacitiesx. This might be good news: top builders having similar abilities means they can meaningfully compete, which is positive for the decentralization of the builder market. Second, the inequality worsens as the MEV tier increases. One possible reason is that providers may give significant MEV opportunities to select builders, as we will discuss in Sec. 5.3.
Figure 6: Violin graphs for the distribution of the QCD across three MEV tiers, for top, middle, and tail builders. The QCD measures the inequality of block-build capacity across builders. This figure reveals that such inequality worsens as MEV increases. Within a given MEV tier, inequality is greatest among tail builders, followed by middle builders, and is least among top builders.
Finding II: The inequality in block building is lowest among top builders, higher among middle builders, and most pronounced among tail builders, and the inequality worsens as a slot’s MEV increases.
Loss from inequality.   For each auction, the loss from inequality is the difference between the highest and second-highest true values, reflecting the potential loss a validator may face due to unequal block-building capacities, as a competitive auction only ensures the winning bid slightly exceeds the second-highest true value.
As shown in Tab. 2, the loss from inequality ranges from 5.6% to 11.5% during the periods covered by our dataset. Comparing the two types of loss, we find that inequality is the proposers’ primary source of loss. Note that the time range of the full bids is from April to December 2023, and the inequality in the current builder market may worsen (as discussed in Sec. 5.3).
5 What Caused Inequality in Block-building Capacities
Given the significant losses proposers incur due to the inequality in builders’ block-building capacity, we further investigate the causes of this inequality in this section.
Although the builder market is technically open to all, the barrier to accessing profitable private order flows (Sec. 2.2) forms an implicit access barrier.
To illustrate the importance of private order flow to the success of builders, Fig. 7 plots the fraction of builder income from private order flows. Currently, approximately 80% of the block value comes from private order flows.
Figure 7: Fraction of builder income from private order flows.
Finding III: The private order flow is crucial for builders to win the MEV-Boost auction because it contributes to more than 80% of the block value in over 50% of blocks.
Where can new builders get private order flows? The entities who provide (or sell) private order flows to builders are called private order flow providers. Different providers impose different access barriers to their private order flows. Some providers, like MEV Blocker or MEV-Share, are open to new builders who satisfy certain requirements, while some (e.g., most searchers) are almost opaque to the public.
To understand what causes the access barrier for new builders, we first identify important “pivotal” providers and then analyze the requirements to access private order flows from these representative pivotal providers.
5.1 Pivotal Private Order Flow Providers
Prominent private order flow providers are listed in Tab. 5. Based on their roles in the MEV supply chain (see Fig. 3), we categorize them into three types: searcher, Telegram trading bot, and channel.
We now present our method to quantify the importance of individual providers. An existing approach [53, 54] is to compute the contribution of each provider to the builders’ income (Sec. 2.3) during a specific period. However, this approach only considers the magnitude of impact but neglects the duration (in terms of time or the number of auctions). For example, a single private transaction, such as one3330xf0464b01d962f714eee9d4392b2494524d0e10ce3eb3723873afd1346b8b06e4 transferring 457.55 ETH to the builder as a tip, could contribute significantly to the builder’s income, but this provider may only influence the MEV-Boost auction in a single slot. Thus, we assess the significance of a provider as the number of auctions for which they played an important role; in particular, we focus on the providers that can affect half of the MEV-Boost auctions in over two weeks, thereby capturing their sustained impact.
We define a provider $P$ as pivotal for an instance of MEV-Boost auction if removing $P$’s transactions from the winning block causes the winner to lose the auction. Intuitively, had the winner not accessed $P$’s order flow, the winner might lose. [^4] Note that the winner may or may not actually lose as it may be able to find an alternative pivotal provider, but this does not invalidate the fact that $P$ is pivotal. We identify pivotal providers in a given auction as follows: for each winning bid, we identify the providers of private order flow within the block and subsequently calculate the revised true value excluding each provider’s contribution. If the recalculated true value drops below the next highest bid (from the partial bids dataset Sec. 3.1), the involved private order flow provider is considered pivotal. Note that there can be multiple pivotal providers for an auction.
Figure 8: A provider $P$ is pivotal if the winner of an auction could have lost without order flows from $P$. The pivotal level of top-5 pivotal providers over time. A line may not start from the beginning because the provider did not exist at that point. Maestro became an independent provider around Dec 2023.
For each pivotal provider, we compute the fraction of auctions in which it is pivotal. We call this metric the pivotal level, a number from 0 to 1. Fig. 8 plots the pivotal level of the top-5 providers. A builder who lacks access to any of their order flows will lose the majority of auctions. However, we found that order flows from these providers are inaccessible to all builders, which requires further investigation. Other observations from Fig. 8 are provided in Appx. F.2.
5.2 Trust Barrier
A direct reason for the access barrier is that private order flow providers are at a disadvantage when interacting with builders and must trust builders not to unbundle the transactions. However, in practice, builders occasionally unbundle searcher’s bundles. For example, the f1b builder unbundled a bundle from the searcher jaredfromsubway.eth at block height 18476515. Although the searcher was compensated afterward, this incident highlights the fragile security foundation of the MEV-Boost auction.
Due to the lack of trustworthy fair exchange mechanisms, providers have to rely on an informal reputation system for self-protection, sending order flows only to builders who meet their requirements. To understand the access barrier caused by the trust crisis, we investigate different providers and their requirements for accessing order flows.
Searchers.   Since searchers are operated by anonymous entities without any public communication channels, there is no public statement about which builders may access their flows. Existing studies indicate that searchers tend to send their order flow to the top builders [81] or share it exclusively with vertically integrated builders [81, 80, 40].
Telegram trading bots.   The requirements for Telegram trading bots, in terms of sending order flow to builders, depend on whether they utilize channels like MEV Blocker [56, 48]. If so, their requirements are aligned with those of these channels. Banana Gun does not explicitly state the requirements for builders but we observe its preference for Titan, which we will discuss later.
Channels.   Since channels are public-facing services, we were able to extract their access requirements from the documentation and/or through their communication channels. The requirements of each provider are summarized in Tab. 5.
The channels impose various requirements on a builder’s reputation. Market share—the percentage of blocks built by a given builder—is a metric most relay providers use to evaluate builders’ reputations. Intuitively, a builder with a high market share, seen as reputable, is less likely to compromise private order flows as doing so will forgo their future revenue. MeowRPC [58] sends the order flow to trusted builders with high market share. Similarly, MEV-Share [37] requires that a builder possesses a competitive market share. MEV Blocker [60] requires that a builder must maintain at least 1% market share over each week of one month.
Subsidization.   New builders face a “chicken and egg” problem as they need access to private order flows to win auctions and gain market share, but the private order flow providers only serve builders with a decent market share. In practice, builders pay out of pocket to overbid in MEV-Boost auctions, a practice known as block subsidization [36], representing an entry barrier for new builders. We quantify the minimal cost for a new builder to meet the MEV Blocker’s requirements (Appx. F.5), which reflects the costs of building a reputation. The minimal cost continues to increase and can be up to 7.4 ETH. This cost represents a lower bound since we do not have the corresponding numbers for other providers. Note that although subsidization can help meet the requirements set by channels, it cannot ensure that searchers’ trust concerns are addressed, allowing builders to receive their order flows.
5.3 Incentive for Integration
Another access barrier to private order flows is that providers exclusively share order flows with specific builders, known as integration [40]. We compute each provider’s builder-specific pivotal level to understand the relationship between providers and builders. For a given provider $P$, its pivotal level specific to builder $B$ is the pivotal level of the MEV-Boost auctions where $B$ is the winner.
Figure 9: Weekly builder-specific pivotal level for top-3 builders. MEV-Blocker and MEV-Share are hidden for clarity.
Fig. 9 plots selected providers’ builder-specific pivotal levels for the top three builders, ranked based on their weekly market share in July 2024. We include providers whose weekly pivotal levels exceed 25% for any of the three builders over two months to illustrate these builder-specific pivotal levels. MEV Blocker and MEV-Share are excluded from this analysis, as all builders connected to them reportedly receive the same order flows. We note that some providers have a pivotal level exceeding 50% only with specific builders, suggesting a possible collaborative relationship. The searcher 0xA69b… consistently influences over 50% of the MEV-Boost auctions won by Beaver, while the searchers 0x0087…, 0x2800…, and 0x51c7… successively affect those won by Rsync. These pivotal providers are all integrated searchers of the corresponding builders that have been identified in previous works [42, 40]. This observation corroborates findings from previous studies [42, 40, 80], confirming that the order flows from integrated searchers are crucial to those builders.
Furthermore, we observe a clear variation in the importance of providers for different builders. For example, the significance of Banana Gun to Titan is greater than it is to other builders. We further investigate the relationship between Banana Gun and Titan and suppose that Banana Gun only sends valuable order flow to Titan when MEV is significant. A case in this point is slot 8614161, where only Titan received the valuable order flow from Banana Gun. The much lower bids from other builders suggest they didn’t receive this order flow; otherwise, they could have outbid Titan. We illustrate this new integration pattern in Appx. F.3. The results in Fig. 15 indicate that this integration pattern also applies to jaredfromsubway.eth and Beaver. Both integrations emerged around January 2024.
There is a clear incentive for integration between providers and builders. If MEV from a provider can determine the auction outcome, sending order flow to an integrated builder is better than sending it to multiple builders. The integrated builder only needs to bid slightly higher to win, and they can share the surplus MEV. However, this integration is detrimental to proposers, as it exacerbates the inequality in builders’ block-building capacities and can eventually lead to significant losses for them.
Finding IV: Two providers (Banana Gun and jaredfromsubway.eth) exclusively share high-value order flows with specific builders (Titan and Beaver) respectively, suggesting a new integration pattern.
6 Related Work
Empirical study on Ethereum’s builder market.   Several previous works studied the centralization of the builder market shares. An early exploration by Yang et al. [92] analyzed the market share of relays and builders. Their findings indicated a centralization of Ethereum’s builder market, with Flashbots and builder0x69 accounting for over 53% of the market share between September 15th, 2022, and November 30th, 2022. Similar studies on the builder market conducted by [84, 41] also confirmed the centralization in the builder market. Their observations were based on the market share and the distribution of total profits generated within the PBS ecosystem. Moreover, several online dashboards [87, 49, 7, 36] offer a dynamic view of the ecosystem’s trends by presenting real-time PBS analytics. However, these studies do not quantify the causes of centralization or the security implications of a centralized builder market.
Private order flow.   Many studies also discuss the importance of private order flows, particularly focusing on private order flows from integrated searchers. Thiery [80] counted the total number of transactions, along with their values, from June 1 to July 15, 2023. He found that private order flows account for only 30% of the transactions yet represent 80% of the total value paid to builders. Our study highlights a similar result: in over 50% of blocks, private order flows account for more than 60% of the block value. Gupta et al. [40] confirmed that builders are more likely to win the MEV auction when their integrated searchers can provide high-value exclusive order flow by effectively exploiting arbitrage between a centralized exchange (CEX) and a decentralized exchange (DEX). The dashboard by Winnie [91] and the study by Titan [81] analyzed the vertical integration between searchers and builders through the quantity and value of transactions sent from searchers to builders. Heimbach et al. [42] observed that during their study, Beaver Builder received a total of 1,941.1 ETH in transaction fees but paid proposers 6,620.94 ETH. This large gap suggests that Beaver might have received significant off-chain profits from its integrated searchers. These studies show that three integrated builders (Beaver, Rsync, and Manta) gain a significant advantage in order flow from their integrated searchers. Our study made the same observation and further quantified the pivotal level.
Strategic behaviors in MEV-Boost auctions.   The strategies used by different entities in the MEV-Boost auctions and their potential impacts have attracted significant interest from researchers. The strategic behaviors of proposers were first studied by [76, 84] focusing on the timing game where proposers strategically delay block proposals to maximize their profits. Öz et al. [66] further conducted an agent-based simulation to analyze how waiting games affect consensus stability, finding that a delay strategy can be profitable and not degrade consensus if sufficiently many validators adopt it. A recent study [1] empirically investigated timing games in the wild and how block proposal delays affect the block fork rate. The results indicate that three validator sets–P2P, Kiln, and Attestant, that are known or suspected of playing timing games–did not delay their block proposals enough to cause their blocks to be forked. Wu et al. [90] proposed a game-theoretic model for MEV-Boost auctions, using simulations to examine various builders’ bidding strategies, including naive, adaptive, last-minute, and bluff bidding. Their results highlighted the importance of latency and exclusive order flow in the effectiveness of bidding strategies. Pai et al. [67] analyzed the latency advantage of searcher-builder integration in adjusting bidding values. The strategies discussed in these papers highlight the potential factors that can make MEV-Boost auctions uncompetitive or inefficient, while our paper quantifies these uncompetitive and inefficient auctions.
Causes and implications of centralized builder market.   A provider may exclusively share its order flow with a single builder. Previous works such as [46, 38, 40] pointed out that exclusive order flows have potential centralizing effects on the builder market, because builders with exclusive order flows are more likely to win MEV-Boost auctions, thus gaining further advantages in receiving order flows from non-integrated searchers and in winning order flow auctions. The decentralization of Ethereum’s builder market not only benefits itself but also positively impacts the decentralization of validators. A recent study showed that heterogeneity among validators in building blocks may lead to centralization; however, a decentralized and competitive builder market can encourage validator decentralization [4]. We complement these theoretical analyses with empirical studies to analyze potential problems caused by the current centralized builder market.
Blockchain decentralization metrics.   Nakamoto coefficient was proposed to describe the minimum number of entities required to disrupt the blockchain’s subsystem [77]. Other metrics, such as the Gini coefficient [18], Shannon entropy [51], and Herfindahl–Hirschman index [73], have also been utilized to measure decentralization (or the lack of it) in staking distribution [39, 47], mining power [52, 89, 47], and block building [41]. An SoK paper [93] comprehensively summarized the existing works on blockchain centralization; however, none explored decentralization from the perspective of entry barriers or auctions.
7 Discussion
7.1 Mitigation
Our results suggest that proposers suffer from significant losses because of builder centralization, which can lead to two undesirable outcomes: First, PBS could become unstable. Second, MEV measured by the auctions may be inaccurate.
In this section, we propose a structural change to the existing MEV supply chain Sec. 2.2 and a solution based on the new supply chain structure to address these two problems.
In the new supply chain, only one builder receives order flows from providers, and proposers can only accept blocks from this builder.
Protocol.   Our solution involves having the builder run the truthful mechanism in [57] in a Trusted Execution Environment (TEE) [15] The pseudocode for the TEE builderis described in Fig. 17.
The builder provides two functions. First, it opens RPC endpoints to receive the private order flows from the providers. Providers send their order flows along with a bid indicating the MEV in their order flows. Second, when it receives a signal for block building (e.g., a new slot starts in Ethereum), it will run an ordering algorithm to build a block with optimal MEV based on the received order flows (a reasonably performant ordering algorithm is described in Appx. E).
The TEE builder runs the truthful mechanism in [57]: it refunds all providers for their marginal value contribution to the block. For each included order flow, the builder builds an alternative block without that specific order flow to calculate its marginal value contribution by comparing the MEV extracted between these two blocks. Next, we outline the key advantages of this solution.
Security guarantees.   TEEs offer a secure and isolated environment, ensuring that private order flows remain confidential and protected from external tampering. Moreover, order flow providers can utilize remote attestation to verify that the builder running within the TEE is indeed the intended one, thereby effectively mitigating trust concerns.
Incentive compatibility.   [57] shows that for providers, truthfully bidding is a dominant strategy when they are refunded based on their marginal value contribution. The formal proof and the required assumptions are in [57]. Here, we offer an intuitive explanation: a provider effectively pays the second price for their order flows when refunding their marginal value contribution.
Analysis.   Since truthfully bidding is an incentive-compatible dominant strategy for all providers, the TEE builder can compute the total MEV in a block from the providers’ bids, achieving an accurate MEV oracle. Moreover, because the builder program in the TEE can be programmed to pay the proposer fairly, the instability of PBS is also addressed.
Future works.   We leave the analysis of whether the new supply chain structure is sustainable in the long term for future work. Intuitively, since providers are incentivized to send order flows only to the builder due to refunds, other builders will naturally be driven out of the market, leaving proposers to only accept blocks from the TEE builder. Furthermore, since real-world proposers might behave rationally, exploring the new equilibrium that emerges among a rational provider, builder, and proposer is another intriguing question to be addressed. Finally, this solution provides an accurate MEV oracle and transparent MEV information. These insights open up new possibilities for future MEV mitigation directions.
7.2 Limitations of Our Approach and Data
True value.   We compute the true values of bids using a similar approach as that in MEV analysis tools [20, 31] and previous works [41, 80]. This approach has a few limitations. First, it relies on the knowledge of the set of addresses controlled by each builder. Most builders use well-known addresses to receive payments from searchers and make payments to proposers, but it is possible that some builders use less publicized addresses. We cross-validated our builder address dataset with other sources to ensure completeness (see Sec. 3.1), but if we missed a builder’s address, the computation of extracted value (and hence true value) would be inaccurate. Secondly, this approach neglects off-chain profits, such as off-chain payments received from searchers. Off-chain payments are more likely to happen between integrated builder-searcher pairs. We might underestimate the true value of a block due to unknown off-chain payments, which could lead to the misclassification of an efficient auction as inefficient due to subsidization. However, in our study, we did not observe any excessively large subsidies. Furthermore, underestimating the winner’s true value does not affect our assessment of competitiveness. We might misclassify an uncompetitive auction as competitive if we underestimate the second-highest true value. However, such underestimation should be uncommon, as we observe that most builders bid around our computed true value. It would be irrational for builders not to bid higher to win the auction if they actually have a higher true value.
Missing pivotal providers.   Our results in Sec. 5.1 might omit some pivotal providers since the providers we can identify are a subset of all providers. For example, BloXroute (a widely used channel) does not publish the list of transactions they processed. Thus, we can’t evaluate its pivotal level. This limitation is common in existing studies [53, 54]. However, there will not be false positives (classification of a non-pivotal provider as pivotal) because we identify a given transaction’s provider based either on the official dataset from MEV-Share and MEV Blocker or on the specific searchers’ addresses (see Sec. 3). Thus, the significance of each provider will only be underestimated rather than overestimated.
Our auction dataset might miss partial bids since they are self-reported by relays. Recall that we used the second-highest bid value in an MEV-Boost auction as the threshold to determine a pivotal provider. Therefore, underestimating the second-highest bid value might also lead us to underestimate the significance of each provider.
7.3 Security Problems in PBS
As we mentioned in Sec. 5.2, one reason for the access barrier to private order flows is the difficulty of establishing trust. A malicious builder, relay, or proposer may attack providers. A real-world incident involved a malicious proposer who double-signed a header to unbundle a builder block and further stole MEV from searchers by proposing a different block [61].
Another security concern is censorship. Ethereum’s censorship resistance currently depends heavily on whether builders choose to censor, particularly the actions of the top builders. For example, before the top builder, Beaver, stopped censoring in July 2024, over 50% of blocks were censoring [86]. Our study reveals that builder centralization remains difficult to overcome, potentially exacerbating the censorship problem in Ethereum.
8 Conclusions
In this paper, we interrogate the causes and implications of builder centralization in Ethereum and challenge the community’s belief that it is acceptable. Our empirical study shows that the current blocker market has led to significant profit losses for validators. These significant losses could lead to PBS’s instability and MEV oracles’ inaccuracy and, if left uncontrolled, can lead to validator centralization. We attribute the loss to two factors: insufficient competition and inequality of block-building capacities, with the latter being the dominant reason. We further investigated the trust and incentive issues of the current MEV supply chain that contribute to builder centralization. To address these two problems, we propose a structural change to the current MEV supply chain and a solution based on this new structure.
Acknowledgements
This work is partially supported by an Ethereum Foundation Grant. We thank ultra sound relay for sharing the full bid dataset. We also thank the teams of BloXroute, Merkle, MeowRPC, MEV Blocker, Flashbots, and Blocknative for answering our questions about access requirements. We are grateful for the insightful comments by Prof. Campbell Harvey and the valuable discussions with Thomas Thiery, Barnabé Monnot, Danning Sui, Burak Öz, Data Always, and Georgios Konstantopoulos.
Appendix A Visualization for Distribution
Fig. 10 shows the distribution of the percentage of builders who submit to the ultra sound relay. Fig. 11 shows the CDF of the winning bid values of MEV-Boost auctions.
Figure 10: Distribution of the percentage of builders who submit to the ultra sound relay.
Figure 11: CDF of the winning bid value of auction dataset.
Appendix B Data Collection
Builder public keys.   We collect builder public keys from two data sources. First, builders often disclose their public keys in official documents. For example, Titan builder has disclosed 15 public keys in its official document [10]. Second, some builders leave their unique marks in the blocks’ “extra data” field. For example, Flashbots builder uses the mark ‘‘Illuminate Democratize Distribute’’. Given that such marks are unauthenticated and susceptible to impersonation, we ignore them unless they are used more than 100 times in the bids by the same builder public key.
Transaction sources.   We identified transactions originated from Telegram trading bots, including Maestro [55], Banana Gun [5], and Unibot [82], by searching for transactions that interacted with their official smart contracts.
Two channels, MEV Blocker and MEV-Share publish data [70, 33] about transactions submitted to them. Specifically, transactions submitted to MEV Blocker are publicly available on Dune [70], from which we extracted 22,819,437 transaction hashes that fall within our study’s range. We obtained 20,334,513 transactions submitted to MEV-Share from the public Flashbots dataset [33]. These transaction labels are sourced from their official datasets and we view them as ground truth.
Appendix C List of Searchers and Builders
In Tab. 3, we provide the list of searchers and their addresses mentioned in our study. In Tab. 4, we have counted the top 25 builders in the market, ranked by their market share from April 2023 to December 2023.
Table 3: The list of searchers
Searcher Address
jaredfromsubway.eth 0xae2Fc483527B8EF99EB5D9B44875F005ba1FaE13
Searcher: 0xA69b… 0xA69babEF1cA67A37Ffaf7a485DfFF3382056e78C
Searcher: 0x0087… 0x0087BB802D9C0e343F00510000729031ce00bf27
Searcher: $0$x280027dd00eE0050d3F9d168EFD6B40090009246
Searcher: 0x51c7… 0x51C72848c68a965f66FA7a88855F9f7784502a7F
Table 4: The percentages of blocks and MEV-Boost blocks built by each builder from April to December 2023.
Builder % of total blocks % of MEV-Boost blocks
Beaver 22.53 24.69
Rsync 20.75 22.74
builder0x69 12.65 13.86
Titan 12.35 13.53
Flashbots 12.24 13.42
BloXroute 2.14 2.34
f1b.io 1.66 1.81
Blocknative 1.40 1.53
eth-builder 0.77 0.84
BuildAI 0.73 0.80
Gambit Labs 0.64 0.70
Jetbldr 0.47 0.51
Manta 0.46 0.51
payload.de 0.39 0.42
boba-builder.com 0.27 0.30
Penguinbuild 0.23 0.25
Eden 0.22 0.24
lightspeedbuilder 0.21 0.23
0x3be…436 0.16 0.18
lokibuilder.xyz 0.12 0.13
bobTheBuilder.xyz 0.11 0.13
blockbeelder.com 0.11 0.12
finest artisanal blocks 0.09 0.10
It’s Free Real Estate 0.08 0.09
nfactorial.xyz 0.06 0.07
Appendix D Implications of Builder Centralization
D.1 Efficiency of MEV-Boost Auctions
Analysis of inefficiency.   Fig. 4 shows that 24.9% of MEV-Boost auctions were not efficient. In an inefficient auction, the builder with the highest true value does not win the auction. There are two possible reasons. First, a builder can win the auction without possessing the highest true value by overbidding, known as block subsidization [36]. Second, builders may “shade” their bids (bid shading means strategically placing bids slightly below one’s true value); if the builder with the highest true value “shades too much,” it will lose.
We analyzed the underlying reasons for all inefficient auctions. In 61.0% of cases, winners win the auctions due to subsidization. In 48.0% of cases, the builder with the highest true value shaded too much. It is important to note that auctions may exhibit inefficiency for multiple reasons. For instance, the winner may subsidize her bid while simultaneously, the builder with the highest true value shades. Thus, the cumulative percentage exceeds 100%.
It is important to note that inefficiency does not necessarily disadvantage the validators. In fact, validators may receive more revenues in inefficient scenarios involving subsidization.
MEV’s implication on efficiency.   The ratio of efficient auctions in each tier of MEV is shown in Fig. 5. We note that the trend of efficiency is less expected. First, only 56.6% of the auctions within the low MEV tier are efficient. This can be attributed to the generally low overall bid values, rendering the auction outcomes vulnerable to alterations from minor subsidies. Second, while auctions tend to be more efficient as MEV increases, we still observe that 77.1% of auctions in the high MEV tier are efficient, and about 86% of inefficient cases result from bid shading, where the builders with the highest true values place their bids much lower than their true values.
D.2 Measuring Block-building Capacities
Coefficient of variation.   We use $\mathsf{CV}(\vec{v})\to\mathbb{R}_{\geq 0}$ to denote the coefficient of variation (CV), a real value in $\mathbb{R}_{\geq 0}$, for values $\vec{v}=(v_{1},\dots,v_{n})$. Specifically, suppose the true values of different builders in slot $s$ are $\vec{tv}_{s}$. The CV of slot $s$ is computed as $CV(\vec{tv}_{s})$.
Figure 12: Violin graphs for the distribution of the CV across three MEV tiers, for top, middle, and tail builders. The CV measures the inequality of block-build capacity across builders.
As shown in Fig. 12, we can make the same finding as Fig. 6. It reveals that such inequality worsens as MEV increases. Within a given MEV tier, inequality is greatest among tail builders, followed by middle builders, and is least among top builders.
D.3 Implications of Uncompetitive Auctions
Fig. 13 shows the distribution of the competitive index (CI) of the MEV-Boost auctions in our study.
Figure 13: The distribution of the CIs of MEV-Boost auctions
Appendix E Reference Builder Algorithm
Flashbots has open-sourced a Rust implementation of the MEV-Boost builder [34], which serves as a reasonably performant reference for builder algorithms. We will briefly introduce their approach. The builder processes each bundle (where a single transaction is considered a bundle containing one transaction) by computing its effective gas price and total profit. Subsequently, the builder sorts all received bundles in descending order based on either effective gas price or total profit. Employing a greedy algorithm, the builder then appends bundles from this sorted list. 1 presents the pseudocode of their algorithm.
Listing 1: Flashbots block-building algorithm
# Compute the effective gas price and profit for a given bundle
# A single transaction is recognized as a bundle of size 1
def receive(bundle):
effective_gas_price, profit = compute_price_and_profit(bundle)
bundle.egp = effective_gas_price
bundle.profit = profit
return bundle
# Building a block
def build(bundles, sorting_key):
# Sort bundles by price or profit
if sorting_key == ’egp’:
bundles.sort(key=lambda b: b.egp)
else:
bundles.sort(key=lambda b: b.profit)
# Initialize
B = []
# Get the current state
state = get_current_state()
# Greedy algorithm
while not bundles.is_empty():
# Get the first bundle from the queue
bundle = bundles.pop(0)
# Simulate the execution of the bundle
success = simulate_bundle(state, bundle)
# Update with the successful bundle
if success:
B.append(bundle)
state.update(bundle)
# Compute payment based on the final state
payment = compute_payment(state)
# Append the payment to the block
B.append(payment)
# Return the valid block
return B
Appendix F What Caused Inequality in Block-building Capacities
F.1 Importance of Private Order Flows
To quantify the importance of private order flows for building valuable blocks, we compute the ratio of block values derived from private transactions. We first identify all private order flows using methods described in Sec. 3.1. Then, for each block, we compute the ratio between values extracted from private transactions (i.e., the total transaction priority fees [21] and direct transfers [32] to the block builder) to the total values extracted from all transactions in the block.
F.2 Pivotal Private Order Flow Providers
We can make a few observations from Fig. 8. First, a particular searcher, jaredfromsubway.eth, has been gaining prominence since February 2023 and reached its peak significance in May 2023. During this period, the winners of more than 70% of the MEV-Boost auctions would have failed had they not received order flow from jaredfromsubway.eth. Its influence has gradually decreased since then but still remains pivotal in 20%-40% of auctions since Nov 2023. Besides, the order flow from Telegram trading bots has become vital for builders with the rise of such bots. For instance, Banana Gun is pivotal in about 40% of the MEV-Boost auctions in September 2023.
Second, we can observe a positive trend from Fig. 8 that MEV Share and MEV Blocker have been increasingly pivotal since May 2023. These two providers positively impact decentralization because they have relatively clear requirements for accessing their order flows (the requirements of different providers will be discussed in Sec. 5.2).
Third, we observe a notable downtrend in the pivotal level of MEV-Share and MEV Blocker starting from November 2023, accompanied by an uptrend in Maestro, a popular Telegram trading bot [55], beginning in December 2023. This is because Maestro used to be an important upstream of MEV-Share and MEV Blocker but stopped in November 2023, causing the pivotal levels of MEV-Share and MEV Blocker to drop significantly.
Although platforms like MEV-Share and MEV Blocker positively affect decentralization, the incentives of their upstream providers may not be aligned. For example, Mastro may prefer to send flows to builders directly for a better price than what MEV-Share and MEV Blocker can offer. It is surprising to see that upstream providers can significantly swing the significance of platforms.
Finding V: We identified five pivotal providers (MEV-Share, MEV Blocker, jaredfromsubway.eth, Banana Gun, Maestro) who had a sustained influence on the outcome of more than half of MEV-Boost auctions. A positive trend towards decentralization is observed as the pivotal level of the searcher jaredfromsubway.eth decreased, while the pivotal level of MEV-Share and MEV Blocker increased. However, their reliance on upstream providers like Maestro is alarming.
Case study for maestro.    To delve deeper into the relationship between these two trends, we plot Maestro’s contribution to the total daily value of transactions sent through MEV-Share and MEV Blocker during our study in Fig. 14. Surprisingly, during the time periods when MEV-Share and MEV Blocker were highly pivotal in Fig. 8, i.e., from May to November, Maestro’s contribution accounts for approximately 80%! This means Maestro used to be an important upstream of MEV-Share and MEV Blocker. In November 2023, Maestro stopped sending transactions to MEV-Share and MEV Blocker, causing the pivotal levels of MEV-Share and MEV Blocker to drop significantly.
Figure 14: Daily percentage of the value of the transactions originating from Maestro for both MEV-Share and MEV Blocker.
Figure 15: Weekly builder-specific pivotal levels for the top-3 builders for slots with winning bid values of no less than 0.5 ETH.
F.3 New Integration Pattern
To illustrate the new integration pattern between providers and builders, we consider the MEV significant in a slot if its winning bid value is no less than 0.5 ETH. Additionally, we focus only on the provider that is the sole pivotal provider in a slot, meaning this provider provides significant MEV to the builder, enabling the builder to win the auction. We apply the above adjustments to builder-specific pivotal levels, and the results are shown in Fig. 15. Compared to Fig. 9, we observe that Banana Gun and Titan, as well as jaredfromsubway.eth and Beaver, follow this integration pattern: a provider selectively shares valuable order flows with specific builders when MEV is significant.
F.4 Requirements of Providers
Tab. 5 summarizes the providers we identify and their requirements for accessing their private order flows.
Type Provider Requirements
Searcher Individual searchers Requirements are unclear.
Telegram trading bot Banana Gun [5], Maestro [55], Unibot [82] The requirements are the same as the channels they use or unclear.
Channel BloXroute [9] Builders are granted access to different tiers, each providing varying privileges and permissions, based on an assessment.
Merkle [59] Builders must have a proven track record of accurately and honestly computing bundles, or they should consult with the Merkle team for consideration.
MeowRPC [58] The transactions are sent to the top builders which MeowRPC’s team trusts.
MEV Blocker [60] Builders must maintain at least 1% market share each week of the previous month.
MEV-Share [37] Builders must demonstrate a competitive market share for consideration.
Transaction Boost [8] Consideration for partnership is based on the Blocknative team’s evaluation of the builder’s reputation.
  • $\star$
    We do not claim that the list of the private order flow providers is complete since there is no ground truth.
Table 5: Collections of the private order flow providers and their requirements.
F.5 Minimal Cost for 1% Market Share
We compute the minimal subsidy needed to establish a 1% market share for one month (which is the requirement of MEV Blocker [60]) by replaying historical auctions with an added new builder who can only access the public order flow. We assume that all builders have access to public order flows but the new builder cannot access any of the private order flows. Furthermore, this new builder can win the auction if their bid value is the same as other builders. Other builders will not submit bids beyond their true value, defined as the total revenue of their blocks. Or they will not increase subsidies if they already subsidize. For simplicity of computation, we assume that one month has 30 days.
Under our assumptions, a new builder must submit a bid at least equal to the auction winner’s maximum–the higher of its bid value or true value. The revenue from the new builder’s block equals the value contributed by the public order flows to the winning bid. Thus, the subsidy needed for winning the auction is the difference between the revenue from public order flows and the required bid value. Then, as we know the subsidy for proposing one block, we can further calculate the minimum overall subsidy a new builder would need to cover to propose 72 blocks each day over one week (Ethereum produces 7200 blocks a day.)
The result is plotted in Fig. 16. Initially, this cost was nearly zero before May 2023, but it has steadily increased, surpassing 7.4 ETH by July 2024.It is important to recognize that the costs derived from this method only represent the lower bound. Predicting the minimum cost of winning a MEV-Boost auction is challenging, so builders need to pay more to ensure victory. Furthermore, builders may subsidize more blocks than required, as it is also difficult to precisely achieve a 1% market share. In reality, the subsidies new builders need to provide are substantially higher. For example, tbuilder [78], a new builder in the builder market, subsidized 6.8 ETH in total for 645 blocks in the week before March 21, 2024 [36].
Figure 16: Subsidy required to establish a 1% market share, as required by MEV Blocker. This quantifies the cost of accessing private order flows.
Appendix G Pseudocode of TEE Builder
The pseudocode for the builder running in TEEs with a refund mechanism is shown in Fig. 17.
[…]
$orders\leftarrow[]$ // Pending transactions.
$sources\leftarrow\{\}$ // Record the order flow providers.
// Receive order flows from providers.
[…]
$orders\leftarrow orders+[order]$
$sources[order]\leftarrow provider$
// Build best block and refund providers.
On building:
$v,B\leftarrow\textsc{build}(orders)$
$payments\leftarrow[]$
// For each provider, refund its marginal contribution,
 For every $order$ in $B$:
  $provider\leftarrow sources[order]$
  $v^{\prime},\_\leftarrow\textsc{build}(orders-[order])$
  $payment\leftarrow\textsc{transfer}(provider,v-v^{\prime})$
  $payments\leftarrow payments+[payment]$
// Remove all transactions included on chain.
 For every $order$ in $B$:
  $orders\leftarrow orders-[order]$
  Delete $sources[order]$
$B\leftarrow B+payments$
[…]
Figure 17: The pseudocode of TEE builder.